### v4

# Cyclotomic Factors and LRS-Degeneracy

Abstract
We present three new, practical algorithms for polynomials in ${\mathbb{Z}}[x]$: one to test if a polynomial is cyclotomic, one to determine which cyclotomic polynomials are factors, and one to determine whether the given polynomial is LRS-degenerate. A polynomial is LRS-degenerate iff it has two distinct roots $\alpha,\beta$ such that $\beta=\zeta\alpha$ for some root of unity $\zeta$. All three algorithms are based on “intelligent brute force”. The first two produce the indexes of the cyclotomic polynomials; the third produces a list of degeneracy orders. The algorithms are implemented in CoCoALib.
keywords:
Linear recurrence sequence , root of unity , cyclotomic polynomial
2020 MSC:
11R18 , 11B37 , 11C08
journal: arXiv
[…]
[TUK]organization=Fachbereich Mathematik, RPTU University Kaiserslautern–Landau, organizationsep= , addressline=Gottlieb-Daimler-Strasse, city=Kaiserslautern, postcode=D-67663, country=Germany
[…]
[UPA]organization=Fakultät für Informatik und Mathematik, Universität Passau, addressline=Innstrasse, city=Passau, postcode=D-94032, country=Germany
1 Introduction
We present three distinct but related algorithms. Cyclotomic polynomials enjoy many special properties, and it is useful being able to identify them quickly — our first algorithm does this, improving upon the methods in [7]. Our second algorithm finds which cyclotomic polynomials appear as factors: this refines the method of Beukers & Smyth [14]. The third algorithm is an improved method for detecting LRS-degenerate polynomials compared to those given in [9]: recall that a polynomial is LRS-degenerate iff it has two distinct roots whose ratio is a root of unity. In [9] the property was named simply degenerate, derived from the notion of degenerate Linear Recurrence Sequence (LRS).
Interest in LRS-degenerate polynomials derives mostly from the Skolem-Mahler-Lech theorem [13], which characterises the zeroes of a LRS: in particular, it states that if the LRS has infinitely many zeroes then the associated polynomial is LRS-degenerate.
All three of our algorithms are based on “brute force applied intelligently” in contrast to the theoretical elegance of earlier methods presented in [7, 14, 9]. In particular our algorithm for detecting LRS-degeneracy is a realization of the modular approach hinted at towards the end of Section 3 in [9].
Our emphasis is on practical utility, so we do not present asymptotic complexity analyses of the algorithms, since they are of little utility in this context (and would detract from the focus of this paper). We do present timings; in particular, we show that our LRS-degeneracy algorithm is significantly faster than the two which were presented in [9].
Acknowledgements
Abbott began working on this topic while at the University of Passau; he then transferred to Rheinland-Pfälzische Technische Universität in Kaiserslautern, where he is supported by the Deutsche Forschungsgemeinschaft, specifically via Project-ID 286237555 – TRR 195. Mexis is working as part of the Project “CySeReS-KMU: Cyber Security and Resilience in Supply Chains with focus on SMEs” (project number BA0100016), co-funded by the European Union through INTERREG VI-A Bayern-Österreich 2021–2027.
The algorithms presented here have been implemented in C++ as part of CoCoALib [1], from version 0.99856. The timings reported in Sections 3.5, 4.2 and 5.11 were obtained on an AMD Ryzen 9 7900X processor, and are given in seconds.
2 Notation, Terminology, Preliminaries
Here we introduce the notation and terminology we shall use.
We shall often restrict attention to square-free, content-free polynomials in ${\mathbb{Z}}[x]$: a general non-zero polynomial in ${\mathbb{Q}}[x]$ may readily be rescaled by clearing denominators and dividing out the content to yield a content-free polynomial in ${\mathbb{Z}}[x]$. We can easily compute the radical (i.e. product of the factors in a square-free factorization) as $f/\gcd(f,f^{\prime})$ where $f^{\prime}$ denotes the formal derivative of $f$. We shall write the reverse of a non-zero polynomial as $\mathop{\rm rev}(f)\,=\,x^{d}\,f(\frac{1}{x})$ where $d=\deg(f)$.
Definition 2.1.
Let $K$ be a field, and let $\zeta\in K$. We say that $\zeta$ is a root of unity if there is $n\in{\mathbb{N}}_{>0}$ such that $\zeta^{n}=1$. Additionally, we say that $\zeta$ is a primitive $\bm{k}$-th root of unity if $\zeta^{k}=1$ and there is no positive $\kappa<k$ with $\zeta^{\kappa}=1$.
Example 2.2.
Let $k\in{\mathbb{N}}_{\geq 2}$, and let $p=1+ks$ be prime (for some $s\in{\mathbb{N}}$). Let $g$ be a primitive root modulo $p$, then $\zeta=g^{(p-1)/k}=g^{s}$ is a primitive $k$-th root of unity in the finite field ${\mathbb{F}}_{p}$.
We recall below some fairly standard notation.
Definition 2.3.
Let $f=\sum_{j=0}^{d}a_{j}x^{j}\in{\mathbb{C}}[x]$. We define the height of $f$ to be $H(f)\;=\;\max\{|a_{0}|,|a_{1}|,\ldots,|a_{d}|\}$ with the special case $H(0)=0$.
Definition 2.4.
Let $\phi$ denote Euler’s totient function for positive integers: namely for every prime $p$ we have $\phi(p^{r})=p^{r-1}(p-1)$, and if $m,n$ are coprime integers then $\phi(mn)=\phi(m)\,\phi(n)$. In particular, $\phi(1)=1$.
Definition 2.5.
Let $\mu$ denote the Möbius function for positive integers: namely, if $n$ is square-free with $r$ prime factors then $\mu(n)=(-1)^{r}$, otherwise $\mu(n)=0$. In particular, $\mu(1)=(-1)^{0}=1$.
Definition 2.6.
Let $k\in{\mathbb{N}}_{\geq 1}$. We write $\Phi_{k}$ to denote the $\bm{k}$-th cyclotomic polynomial (or cyclotomic polynomial of index $\bm{k}$), i.e. the unique monic polynomial of smallest degree whose roots are all primitive $k$-th roots of unity. It is an elementary result that $\Phi_{k}(x)\in{\mathbb{Z}}[x]$ and $\deg(\Phi_{k})=\phi(k)$.
Lemma 2.7.
We recall two product relations for cyclotomic polynomials:
  • (a)
    The classical product
    $x^{k}-1\;=\;\prod_{d\mid k}\Phi_{d}(x)$
  • (b)
    Applying Möbius inversion to (a) gives
    $\Phi_{k}(x)\;=\;\prod_{d\mid k}(x^{d}-1)^{\mu(k/d)}$
Definition 2.8.
For a positive integer $k$ we write $G_{k}$ to denote the $\bm{k}$-th Graeffe transformation [17] which maps roots to their $k$-th powers. Specifically we define $G_{k}(f(x))\,=\,\mathop{\rm res}_{y}(f(y),\,x-y^{k})$.
Remark 2.9.
The Graeffe transform exists in the same ring as the original polynomial. Also for $k=2$ and $k=3$ the Graeffe transform can be computed quickly and easily: e.g. $G_{2}(f)\,=\,(-1)^{d}(f_{even}^{2}-x\,f_{odd}^{2})$ where $d=\deg(f)$, and $f_{even}$ and $f_{odd}$ are defined by $f(x)\,=\,f_{even}(x^{2})+x\,f_{odd}(x^{2})$.
For composite $k$ we can compute a succession of simpler Graeffe transforms based on $G_{ab}(f)\,=\,G_{a}(G_{b}(f))$.
If several Graeffe transforms for consecutive $k$ are needed, one can use one of the approaches mentioned in [11]; we implemented approach 4 from that work, based on Newton symmetric functions, due to it being iteratively applicable to consecutive orders $k$.
Definition 2.10.
Let $k\in{\mathbb{N}}_{\geq 2}$ and non-zero $f\in{\mathbb{C}}[x]$. Then we say that $f$ is $\bm{k}$-LRS-degenerate iff $f$ has two distinct roots $\alpha,\beta\in{\mathbb{C}}$ such that $\alpha/\beta$ is a primitive $k$-th root of unity. We also say simply that $f$ is LRS-degenerate to mean that there exists at least one order $k\in{\mathbb{N}}_{\geq 2}$ such that $f$ is $k$-LRS-degenerate. We are primarily interested in the case $f\in{\mathbb{Z}}[x]$.
Definition 2.11.
Let non-zero $f,g\in{\mathbb{C}}[x]$. Then we say that $f$ and $g$ are LRS-degeneracy equivalent iff there are scale factors $\lambda,\mu\in{\mathbb{C}}_{\neq 0}$ such that $f(x)=\mu\,x^{d}\,g(\lambda x)\;$ or $f(x)=\mu\,x^{d}\,g(\lambda/x)$ for some $d\in{\mathbb{Z}}$. It is elementary to show that this is an equivalence relation.
Remark 2.12.
If $f$ and $g$ are LRS-degeneracy equivalent then $f$ is $k$-LRS-degenerate iff $g$ is $k$-LRS-degenerate.
3 Algorithms for Testing Cyclotomicity
An elegant method for determining whether a polynomial is cyclotomic was presented in [7]. We present a computational approach which is more effective than elegant; it assumes that the polynomial is represented explicitly, and that we can easily obtain the coefficients of the terms of highest (or lowest) degree.
Let $f\in{\mathbb{Z}}[x]$ be a non-constant, content-free polynomial. We explain how we check whether $f$ is cyclotomic; and if it is indeed cyclotomic, we determine its index.
We start with a few trivial “quick checks” which every cyclotomic polynomial will pass (but also some other polynomials may pass too). We handle specially the case where $f$ is of the form $g(x^{r})$ for some exponent $r>1$: this permits us to concentrate on identifying only cyclotomic polynomials with square-free index. If all the quick checks pass, we apply a more costly absolute test.
3.1 Preliminary quick checks
The first 3 checks are very simple and quick, and should be applied in the order presented.
  • Q1
    If $f$ is not monic, it is not cyclotomic. We henceforth assume that $f$ is monic.
  • Q2
    If $\deg(f)$ is odd, it is not cyclotomic, except for the following two cases: $f(x)=x-1=\Phi_{1}(x)$ or $f(x)=x+1=\Phi_{2}(x)$. We henceforth assume that $\deg(f)$ is even.
  • Q3
    If we can obtain the constant coefficient of $f$ quickly, we can check that this is $1$ — otherwise $f$ is not cyclotomic.
We defer testing palindromicity until later: see check Q5b in Section 3.4.
3.1.1 Coefficient of $x^{d-1}$ and the special case $f(x)=g(x^{r})$
  • Q4a
    Check the coefficient of $x^{d-1}$ where $d=\deg(f)$. If it is not 1-1, 0 or $1$ then $f$ is not cyclotomic. If it is 0, we take special action (described immediately below: Q4b–Q4e).
  • Q4b
    We check two easy special cases. If $f(x)=x^{d}{-}1$ with $d>1$ then it is not cyclotomic. If $f(x)=x^{d}{+}1$ with $d>1$ then it is cyclotomic iff $d=2^{s}$ in which case the index is $2d$.
  • Q4c
[…]
  • Q4d
[…]
  • Q4e
    Otherwise we conduct the “square-free index” test of Section 3.3 on $g(x)=f(x^{1/r})$. If $g(x)$ is not cyclotomic then neither is $f(x)$. If however $g(x)$ is cyclotomic with index $k$, we then check whether $\mathop{\rm rad}(r)\mid k$. If so then $f(x)$ is cyclotomic with index $rk$, otherwise $f$ is not cyclotomic.
Example 3.1.
[…]
In contrast, with $f=x^{6}+x^{3}+1$ we obtain $r=3$ in Q4c, and as before we see that $g=\Phi_{3}$, but this time $\mathop{\rm rad}(r)\mid k$; so we conclude in Q4e that $f=\Phi_{9}$.
3.2 Further quick checks
We present some further simple checks; unlike the checks Q1–Q4b, these checks can access all the coefficients of the polynomial, so they are typically costlier. Checks Q5c and Q5d are valid only for cyclotomic polynomials with square-free index, so they must be conducted only after all checks in Section 3.1 have been made; in contrast, checks Q5a and Q5b are independent of the other checks.
  • Q5a
    Check that $f$ is palindromic.
  • Q5b
    Check that the coefficient height of $f$, denoted ${\rm H}(f)$, is within bounds: here we can combine information from sequence A138474 at OEIS [16], and from a table of height bounds based on degree (see also Section 3.4.4).
  • Q5c
    Compute $V_{1}=f(1)$; this is just the sum of the coefficients. If $V_{1}\neq 1$ and $V_{1}\neq d{+}1$ then $f$ is not cyclotomic. If $V_{1}=d{+}1$: check that d+1d{+}1 is prime and that ${\rm H}(f)=1$; if so, $f$ is cyclotomic with (prime) index d+1d{+}1; otherwise $f$ is not cyclotomic.
  • Q5d
    Compute $V_{-1}=f(-1)$; this is an “alternating sum” of the coefficients. If $V_{-1}\neq 1$ and $V_{-1}\neq d{+}1$ then $f$ is not cyclotomic. If $V_{-1}=d{+}1$: check that d+1d{+}1 is prime and that ${\rm H}(f)=1$; if so, $f$ is cyclotomic with index 2(d+1)2(d{+}1); otherwise $f$ is not cyclotomic.
Observe that all four of these checks can be performed together in a single scan over the coefficients. Also, checks Q5c and Q5d can be seen as a special case of our method for identifying cyclotomic factors presented in Section 4.
Remark 3.2.
We implemented these checks in CoCoALib, but found that they are not as fast as the “matching prefix” method described in Section 3.3. This is partly explained by the fact that in CoCoALib “reading” all the coefficients of a polynomial is costly. Thus our final implementation does not actually employ these checks.
3.3 Testing for a cyclotomic polynomial with square-free index
Here we assume that all the checks from Section 3.1 have been made, so that if $f$ is cyclotomic then its index must be square-free (and at least 3). As already indicated, the test we give here involves “smart brute force”, and is decidedly more computationally costly than the quick checks in Section 3.1.
The idea here is simple: we generate a list of possible candidate indexes, and then repeatedly whittle it down by computing prefixes of the corresponding cyclotomics, and keeping only those indexes whose prefixes match the highest terms in $f$. If more than one candidate index survives, we refine the list again using longer prefixes. In Algorithm 1 (CycloIndex) we write $f_{m}$ to mean the $m$-prefix of $f$, i.e. all terms of degree greater than dmd{-}m.
Remark 3.3.
In Algorithm 1 (CycloIndex) in Steps 16 and 17, it suffices to verify the $\frac{d}{2}$-prefix and then check that $f$ is palindromic.
Algorithm 1 (CycloIndex)
[…]
2:
[…]
4:Filter $L$: keep only those $n\in L$ for which $\mu(n)=-a_{d-1}$
5:Choose initial prefix length $m=32$
6:while $L$ contains more than 1 element do
7:  $L_{new}\leftarrow[\,]$, an empty list
8:  for each $n\in L$ do
9:   Compute $p$, the $m$-prefix of $\Phi_{n}$
10:   if $p=f_{m}$ then append $n$ to $L_{new}$ end if
11:  end for each
12:  $L\leftarrow L_{new}$
13:  $m\leftarrow 4m$
14:end while
15:if L=[]L=[\,] then return not cyclotomic end if
[…]
17:if $f=\Phi_{k}$ then return $k$ else return not cyclotomic end if
Remark 3.4.
In Algorithm 1 (CycloIndex) the initial value for the prefix length $m$, and the increment in Step 13 were chosen because they worked well in our tests. We did not try “optimizing” the strategy as it seemed to be unnecessary. After the first iteration of the main loop, $L$ typically contains only very few elements.
Remark 3.5.
A referee described a more efficient version of Algorithm 1 (CycloIndex): an outer loop iterates over the candidate indexes, and an inner loop checks each candidate by constructing successively longer prefixes until the candidate is either excluded or confirmed as sole candidate for final verification. This approach will be faster if the input polynomial is indeed cyclotomic, and its index appears early in the list of candidates. The difference in overall speed is likely minor since the final verification is typically the main part of the total computational cost. Moreover, Algorithm 1 (CycloIndex) as presented can be trivially modified to skip the final verification (Steps 16 and 17): it then determines the index when the input is known to be cyclotomic — see Section 3.5. It is not clear whether the referee’s variant could easily be adapted to forego the final verification.
3.3.1 Computing an $m$-prefix
Let $n\in{\mathbb{Z}}_{>2}$ be square-free. Since cyclotomic polynomials are palindromic, prefixes and suffixes are equivalent. It is convenient here to discuss suffixes: namely $\Phi_{n}(x)\bmod x^{m}$. We can compute an $m$-suffix cheaply via the Möbius inverted product:
$\Phi_{n}(x)\;=\;\prod_{d\mid n}(x^{d}-1)^{\mu(n/d)}\;=\;\prod_{d\mid n}(1-x^{d})^{\mu(n/d)}$
We may negate the factors because $n$ has an even number of divisors. Observe that multiplying by $1-x^{d}$ is just a shift-and-subtract operation, and indeed dividing by $1-x^{d}$ can also be implemented as a shift-and-add operation (working up from lowest to highest degree). We may ignore all $d\geq m$.
3.3.2 Coefficient Bounds and Machine Integers
In Algorithm 1 (CycloIndex) at Step 10 we use $f_{m}$, the $m$-prefix of $f$. We incorporate a coefficient height check while generating $f_{m}$ from $f$; if a coefficient is too large then we know that $f$ cannot be cyclotomic. Furthermore, so long as $m$ is not too large, all arithmetic can be done with machine integers: the outermost coefficients of a cyclotomic polynomial cannot be too large. A table of bounds for each of the outermost coefficients is available as sequence A138474 at OEIS [16]. On a 32-bit computer overflow cannot occur with $m\leq 808$; and on a 64-bit computer all $m\leq 1\,000$ are safe — the table at OEIS stops at index 1 0001\,000; the true limit for 64-bit computers is likely considerably higher.
For prefixes of length m>1 000m>1\,000 we can use the tabulated values of height bounds for cyclotomic polynomials (up to index 2 583 303 5552\,583\,303\,555 — see Section 3.4.4). Additionally, for each candidate index we can apply the formulas below to obtain potentially better bounds (especially when the index has few prime factors). Our implementation gives an error if the best bound found exceeds the largest machine integer.
In [8] there are formulas bounding the height of $\Phi_{n}$ and of its power-series inverse $1/\Phi_{n}(x)$ for an odd, square-free index $n=p_{1}p_{2}\cdots p_{k}$ where $p_{1}<p_{2}<\cdots<p_{k}$. Interestingly, all the bounds are independent of the largest two prime factors. If $k<3$ then the height bound is $1$. If $k=3$ then the bound is ${\textstyle\frac{2}{3}}p_{1}$ according to preprint [12], and on the basis of this result the article [8] establishes bounds on the height of $\Phi_{n}$ for $k=4,5,6$:
$\frac{2}{3}p_{1}^{3}p_{2}\qquad\frac{2}{9}p_{1}^{7}p_{2}^{3}p_{3}\qquad\frac{32}{729}p_{1}^{15}p_{2}^{7}p_{3}^{3}p_{4}$
The bounds on the height of the inverse have a similar form but with different constant factors. If the proof in [12] is not confirmed as correct then [8] obtains other constant factors. A fully general formula is given in [6]. So far we have implemented just the simple cases $k=1$ and $k=2$.
If some of the prime factors of $n$ are greater than the prefix length $m$, we may sometimes obtain lower bounds. For instance, if $n=p_{1}p_{2}p_{3}$ and $p_{1}<m\leq p_{2}$ then the only non-trivial truncated factors in the Möbius product are the same as those required to compute the $m$-prefix of $\Phi_{p_{1}}$. Thus the height bound for index $p_{1}$ limits the $m$ outermost coefficients also of $\Phi_{n}$; indeed, in this instance we know that the coefficients of $x^{e}$ for $p_{1}\leq e<p_{2}$ must be zero. Depending on whether the number of prime factors bigger than $m$ is even or odd, we use the height bound for the cyclotomic polynomial or for its inverse.
3.3.3 Computing preimages under $\phi$
All elements of $\phi^{-1}(d)$ can be computed quickly enough by a simple recursive “tree search” [10]. The algorithm may easily be adapted to produce only the square-free preimages. The algorithm could also produce the preimages naturally in factored form, which is ideal for our application — however we have not yet implemented this “optimization”.
3.4 Testing via evaluation
We recall the quick checks Q5c and Q5d from Section 3.2: these perform evaluation at $1$ and at 1-1 respectively, combined with a check that the resulting value is permissible. We now show that evaluation at $2$ provides us with a simple way of obtaining a unique candidate index (or proof that the polynomial is not cyclotomic). However, our implementation of the prefix method of Algorithm 1 (CycloIndex) turned out to be faster.
3.4.1 Obtaining a candidate index (via remainders of powers)
An irreducible, monic polynomial is cyclotomic iff it divides $x^{k}-1$ for some exponent $k>0$. The smallest such $k$ is equal to the index of the cyclotomic polynomial. We could use this as an absolute test for cyclotomicity: check whether $f(x)$ divides $x^{k}-1$ for $k=1,2,3,\ldots$ Bounds on the inverse totient function ensure that only finitely many values of $k$ need to be checked. This approach was presented in [7]. However, actually doing this is rather inefficient.
We could make the check faster by computing the canonical remainder of $x^{r}\bmod f(x)$ for $r=2d,\,3d,\,4d,\,\ldots$. Bounds on the inverse totient function show that for degrees <36 495 360<36{\,}495{\,}360 (see sequence A355667 at OEIS [16]) we can stop at $r=6d$ since $k$ cannot exceed this value. If the polynomial is indeed $\Phi_{k}$ then once we try the first exponent $r\geq k$ we obtain remainder $x^{s}$ (with s=rks=r-k), and it is easy to identify if the remainder has this form. From this remainder we easily obtain $k$ as rsr-s. However again, actually computing these polynomial remainders would still be rather costly.
Instead we mimic the above but under the mapping $x\mapsto 2$. We evaluate $f(2)$, and check that $2^{d-1}<f(2)<2^{d+1}$. Then we compute $2^{r}\bmod f(2)$ for $r=2d,\,3d,\,4d,\,\ldots,7d$; to advance to the next value in the sequence we just multiply by $2^{d}$ then reduce modulo $f(2)$. Observe that on a binary computer it is easy to detect if the remainder is of the form $2^{s}$, in which case we obtain the candidate index k=rsk=r-s. We finally verify that $\deg(f)=\phi(k)$ and if so, also that $f=\Phi_{k}$ by computing $\Phi_{k}$.
Remark 3.6.
We note that the final check (that $f=\Phi_{k}$) is necessary, at least for higher degrees and index $k$ not of the form $p$ or $2p$ for some prime $p$. Consider the polynomial $\Phi_{k}(x)+x^{d/2-3}\,(x^{2}-1)^{2}\,(2x^{2}-5x+2)$ where $d=\deg(\Phi_{k})$; it is palindromic, and will pass all the evaluation tests, since the added polynomial vanishes at $x=\pm 1$ and $x=2$. For higher degrees the coefficient height check will pass also.
3.4.2 Explicit Algorithm
The algorithm has to exclude just a few edge cases: namely, $\Phi_{1}$ and $\Phi_{6}$ which are exceptions in Zsigmondy’s theorem. It is convenient, but not necessary, to exclude $\Phi_{2}$ as well. We assume that the $\bmod$ function returns the least non-negative remainder.
We note that on a binary computer it is quick and simple to recognize if an integer is a power of $2$ and to determine the corresponding exponent: we do this in Algorithm 2 (CycloIndexByEval) Step 8.
Algorithm 2 (CycloIndexByEval)
[…]
2:
[…]
4:if $v\leq 2^{d-1}$ or $v\geq 2^{d+1}$ then return not cyclotomic end if
[…]
6:for $r^{\prime}\leftarrow 2$ to 77 do $\triangleright$ where $r^{\prime}\leftrightarrow{\textstyle\frac{r}{d}}$
7:  $a\leftarrow(ap)\bmod v$
8:  if $a=2^{s}$ for some $s$ then return $r^{\prime}d-s$ end if
9:end for
10:return not cyclotomic
3.4.3 Supporting arguments for evaluation at 2
We justify Algorithm 2 (CycloIndexByEval) presented in Section 3.4.2.
Lemma 3.7.
Let $k\geq 2$, then $2^{d-1}<\Phi_{k}(2)<2^{d+1}$ where $d=\deg(\Phi_{k})$.
Proof.
We just use Theorem 5 from [15] with $b=2$. Since $\Phi_{k}(2)$ is odd, we can change the inequalities to strict ones. ∎
Lemma 3.8.
Let $k\in{\mathbb{N}}_{>0}$ with $k\neq 6$. Let $r\in{\mathbb{N}}_{>0}$ be such that $\Phi_{k}(2)\;|\;2^{r}{-}1$. Then k|rk\;|\;r.
Proof.
We observe that for $r>0$ we have $2^{r}-1\neq 0$. Since $\Phi_{k}(x)$ divides $x^{k}-1$, we have that $\Phi_{k}(2)$ is non-zero and divides $2^{k}-1$. Moreover, since $x^{k}-1$ divides $x^{sk}-1$ for all $s\in{\mathbb{N}}_{>0}$, we immediately have that $\Phi_{k}(2)$ divides $2^{sk}-1$ for all $s\in{\mathbb{N}}_{>0}$.
To show that there are no other exponents $r$ we invoke Zsigmondy’s Theorem (see Section 4.1.1) with $a=2$ and $b=1$. ∎
Remark 3.9.
We explain here why the exponents we use for testing (namely, $2d,3d,4d,\ldots,7d$) are sufficient.
Assuming the polynomial is $\Phi_{k}$ for some square-free index $k$, we know the value of $\phi(k)=\deg(\Phi_{k})$. Successive maxima of $n/\phi(n)$ arise when $n$ reaches a primorial number: the least $n$ where $n>7\phi(n)$ is at about $n\approx 1.3\times 10^{16}$ when $\phi(n)\approx 1.8\times 10^{15}$. A fully general implementation would compute the greatest exponent to test based on the degree of the polynomial being tested.
Now we explain why we can “make jumps” of size $d$. For any $\delta\in{\mathbb{N}}$ we have that $2^{k+\delta}\equiv 2^{\delta}\pmod{\Phi_{k}(2)}$. So if $\delta<\deg(\Phi_{k})$ then by Lemma 3.7 the least non-negative remainder of $2^{k+\delta}$ modulo $\Phi_{k}(2)$ is $2^{\delta}$, i.e. a power of $2$: something we can test quickly on the computer.
Let $s\in{\mathbb{N}}$ be such that $sd<k\leq(s{+}1)d$. So the least non-negative remainder of $2^{sd}\mod\Phi_{k}(2)$ is not a power of two (by Lemma 3.8), but $(s{+}1)d=k+\delta$ for some $\delta$ with $0\leq\delta<d$. Hence $2^{s(d+1)}\equiv 2^{\delta}\pmod{\Phi_{k}(2)}$, and the right hand side is the least non-negative remainder, which we can recognise as a power of 2.
3.4.4 Height bounds for cyclotomic polynomials
In Q5a we check the height of $f$. The outermost coefficients can be checked via A138474 as in Section 3.3.1, but once that table is exhausted we need another way: we use a second table as described here (up to degree 11 612 15911\,612\,159). If the degree of $f$ is too large for this second table, the check is skipped.
Since there are only finitely many cyclotomic polynomials of each degree, there is a well-defined maximum coefficient height (as a function of degree). We do not know a nice formula for this maximum, but use a precomputed table up to some degree limit. For compactness we employ a non-decreasing table of height maxima up to the given degree, though this does weaken the check.
In our implementation the table of maximum coefficient heights is represented as a cascade of if statements. Letting d denote the degree, the first few lines are:
      if (d < 48) return 1;  // Phi(105)
      if (d < 240) return 2; // Phi(385)
      if (d < 576) return 3; // Phi(1365)
      if (d < 768) return 4; // Phi(1785)
      if (d < 1280) return 5; // Phi(2805)
      if (d < 1440) return 6; // Phi(3135)
      if (d < 3840) return 7; // Phi(6545)
      if (d < 5760) return 9; // Phi(15015)
      if (d < 8640) return 23; // Phi(21945)
Note that our table is not the same as sequence A160340 in OEIS [16]: our table comprises successive maximum heights by degree, whereas A160340 lists successive maximum heights by index. The first few entries do coincide, but then they diverge: e.g. at degree $5760$ due to $\Phi_{15015}$, and at many other points.
Later entries in our table were derived from [3] whose data is now available as part of [2] under Library of data on the heights and lengths of cyclotomic polynomials.
3.4.5 Obtaining a candidate index (via table-lookup)
We mention here a quick and easy way of identifying the index of a cyclotomic polynomial, provided it has square-free index and degree $\leq 1000$. We do not use this in our implementation because of its limited range.
Let $S$ be the set of cyclotomic polynomials of degree up to $1000$ and having square-free index. Let the prime $p=57737$. Then the mapping $S\to{\mathbb{F}}_{p}$ sending $f\mapsto f(2)\bmod p$ is 1–1. So we can precompute a table of indexes: e.g. an array whose $r$-th entry is the index $k$ such that $r=\Phi_{k}(2)\mod p$. We fill “empty” array entries with a value which is not a valid index (e.g. 0).
We can use the table as follows: if $\deg(f)\leq 1000$, compute $r=f(2)\bmod p$. If the table entry corresponding to residue $r$ is “empty” then $f$ is surely not cyclotomic. Otherwise, let $k$ be the table entry, and verify that $\phi(k)=\deg(f)$ and that $\mu(k)$ is equal to the coefficient of $x^{d-1}$; if so then maybe check that $2^{k}\equiv 1\pmod{f(2)}$, and finally compute $\Phi_{k}$ then test whether $f=\Phi_{k}$.
Such a table look-up is quick, but has the obvious disadvantages that the table must be precomputed, and that the table works only over a limited degree range.
3.5 Timings
Here we give some sample timings of our implementation of the “prefix method” (from Section 3.3) in CoCoALib. We restrict to the case of identifying the index of a cyclotomic polynomial without verification in Steps 16 and 17, because the verification step involves computing the full polynomial, and this verification dominates. To present meaningful timings we consider only “difficult” cases where a large prefix length is needed, and there are many preimages to consider: for these examples, the smaller factor of each index is a lower bound on the prefix length required.
Table 1: Timings for CyclotomicIndex
Index Factors $\phi^{-1}(d)$ Time
124 525 451124\,525\,451 $8641\times 14411$ 3 3543\,354 $0.125$
120 507 533120\,507\,533 $4481\times 26893$ 6 8616\,861 $0.140$
334 482 719334\,482\,719 $5279\times 63361$ 8 1458\,145 $0.359$
399 083 849399\,083\,849 $3329\times 119881$ 10 81910\,819 $0.406$
4 Algorithms for Detecting Cyclotomic Factors
In this section we present an effective method of determining which cyclotomic polynomials, if any, divide a given polynomial. Once again, our method is “refined brute force”.
An obvious way to determine the cyclotomic factors would be first to compute the irreducible factors, and then use the method of Section 3 to test each factor for being cyclotomic — this is practical only for moderate degrees. This specific task has already been studied and elegantly solved by Beukers and Smyth: in [14] there is a remarkably short and simple algorithm for finding the product of the cyclotomic factors of a square-free polynomial. However, they did not address the matter of identifying each individual cyclotomic factor. We present a new effective algorithm which identifies quickly a list of indexes of cyclotomic factors. It is possible that the list contains a few false positives, though their presence is “unlikely”; the candidate indexes can be tested by trial division.
Let $f\in{\mathbb{Z}}[x]$ be the polynomial whose cyclotomic factors we wish to find. We list here some potential preprocessing steps — they are not required for correctness, yet could make the overall computation faster (but in our implementation there was no benefit):
  • Take the radical: $f\leftarrow f/\gcd(f,f^{\prime})$
  • Extract the palindromic part: $f\leftarrow\gcd(f,\mathop{\rm rev}(f))$
  • Extract the product of the cyclotomic factors using the Beukers & Smyth algorithm — this can be costly if $f$ has high degree
4.1 Finding cyclotomic factors by evaluation
The basis for our approach is evaluation of the polynomial (at several rational points), and Zsigmondy’s theorem (see [18]).
4.1.1 Zsigmondy’s Theorem
[…]
Zsigmondy’s theorem [18] tells us that if we have the sequence $s_{n}=a^{n}-b^{n}$ for coprime integers $a>b>0$ then for every index the element $s_{k}$ has a primitive prime. The only exceptions are:
  • if $a=b+1$ then for $k=1$ we have $s_{k}=a^{1}-b^{1}=1$ which has no prime factors
  • if $a=2$, $b=1$ then for $k=6$ there is no such prime factor
  • if $a+b=2^{r}$ for some $r$ then for $k=2$ there is no such prime factor
We shall refer to the primitive primes for such sequences as Zsigmondy factors. The specific case $a=2,b=1$ was already solved by Bang’s Theorem (see [5, 4]).
4.1.2 Evaluation at a rational
For compactness we introduce a variant of function call notation. Let $f\in{\mathbb{Z}}[x]$ be non-zero of degree $d$, and let $\beta=p/q\,\in{\mathbb{Q}}_{>0}$ with $p$ and $q$ positive and coprime; then we write
$f(\beta)_{num}\;=\;q^{d}\,f(\beta);$
and if $f$ is zero then we define $f(\beta)_{num}=0$. Thus for all $f\in{\mathbb{Z}}[x]$ and all $\beta\in{\mathbb{Q}}_{>0}$ we have $f(\beta)_{num}\in{\mathbb{Z}}$. Also if $\beta\in{\mathbb{Z}}$ then $f(\beta)_{num}=f(\beta)$.
Lemma 4.1.
Let $\beta=p/q\,\in{\mathbb{Q}}_{>1}$ with $p$ and $q$ coprime and positive. Let the sequence $\sigma_{k}=p^{k}-q^{k}$; applying Zsigmondy’s theorem we obtain a Zsigmondy factor $z_{k}$ for each index $k\in{\mathbb{N}}_{>0}$ — with the noted exceptions. Let the sequence $\tau_{k}=\Phi_{k}(\beta)_{num}$. Then the $z_{k}$ are Zsigmondy factors for the $\tau$ sequence.
Proof.
[…]
4.1.3 Algorithm
We present our algorithm in two parts: the inner Algorithm 3 (called RefineCandidates) which refines a list of candidate indexes, and the outer Algorithm 4 (called FindCycloFactors) that calls the inner one repeatedly.
The inner “refinement” algorithm takes 3 inputs: the polynomial $f$, the evaluation point $\beta\in{\mathbb{Q}}_{>1}$, and a list of candidate indexes. It produces a refined list of candidate indexes, which might be empty.
Algorithm 3 (RefineCandidates)
1:Polynomial $f$, evaluation point $\beta=p/q\in{\mathbb{Q}}_{>1}$, and $L$ a list of candidate indexes (excluding 11 and 22)
2:
[…]
5:if $N=0$ then $f\leftarrow f/(qx-p)$; go to Step 3 end if
[…]
7:for each index $k\in L$ do
8:  if $\phi(k)>\deg(f)$ then continue to next $k$ end if
9:  $g\leftarrow\gcd(p^{k}-q^{k},N)$
10:  if $g=1$ then continue to next $k$ end if
[…]
12:  Append $k$ to $L_{new}$
13:  $N\leftarrow N/\gcd(N,g^{\infty})$ $\triangleright$ See Remark 4.2
14:end for each
15:return $L_{new}$
Remark 4.2.
In Algorithm 3 (RefineCandidates) at Step 10 we write $N/\gcd(N,g^{\infty})$ to mean $\lim_{s\to\infty}N/\gcd(N,g^{s})$; it is the largest factor of $N$ coprime to $g$.
The outer algorithm takes just the polynomial $f$ as input; it could also perform the preprocessing steps mentioned earlier. It checks explicitly for the factors $\Phi_{1}$ and $\Phi_{2}$, so that calls to the inner algorithm are not complicated by the exceptions in Zsigmondy’s theorem. It always chooses $2$ as the first evaluation point (for reasons of computational speed); thereafter it picks random small rationals. The randomness serves to preclude the possibility of constructing a small input which consistently causes the algorithm to perform poorly — this is standard practice in randomized algorithms.
Algorithm 4 (FindCycloFactors)
1:Primitive, non-constant polynomial $f\in{\mathbb{Z}}[x]$
2:
3:Optional: Apply preprocessing steps from beginning of Section 4
[…]
7:repeat
8:  Pick a random evaluation point $\beta\in{\mathbb{Q}}_{>1}$
9:  $L_{new}\leftarrow\textbf{RefineCandidates}(f,\beta,L)$ $\triangleright$ see Algorithm 3
10:  if $L_{new}=[\,]$ or $L_{new}=L$ then return $L_{12}\cup L_{new}$ end if
11:  $L\leftarrow L_{new}$, and perform another iteration
12:until forever
Remark 4.3.
If the input polynomial is monic and of even degree, we can perform, before Step 5, an initial cyclotomicity test (from Section 3) as this is very quick. The initial refinement at $\beta=2$ was quite effective in our tests: i.e. only few false positives survived. As written, Algorithm 4 (FindCycloFactors) has difficulty excluding the indexes $3,4$ and $6$; so in Step 10, the first time the condition $L_{new}=L$ is satisfied, we perform one more iteration but using a $\beta$ value chosen so that each of $\Phi_{k}(\beta)_{num}$ for $k=3,4,6$ has a prime factor larger than 10 00010\,000, e.g. suitable $\beta$ values are $\frac{117}{98}$, $\frac{133}{18}$ or $\frac{169}{6}$.
Remark 4.4.
An advantage of evaluating at rationals rather than just at integers is that the evaluations are usually smaller. There are $\Theta(B^{2})$ reduced rationals with numerator and denominator bounded by $B$. Let $\beta$ be such a rational and $f\in{\mathbb{Z}}[x]$ of degree $d$, then $f(\beta)_{num}$ is bounded by $(d{+}1)\,{\rm H}(f)\,B^{d}$. In contrast, we have about $2B$ integers giving evaluations below this bound.
Remark 4.5.
Let $f=\sum_{j=0}^{d}a_{j}x^{j}$ be non-constant. Let $p$ be a prime which divides $a_{0},a_{1},\ldots,a_{r}$ and $a_{d},a_{d-1},\ldots,a_{d-s}$; note we do not need to compute $p$, it suffices to verify that $\gcd(a_{0},a_{1},\ldots,a_{r},\;a_{d},a_{d-1},\ldots,a_{d-s})\neq 1$. Suppose $\Phi_{k}\mid f$, then $\psi_{p}(\Phi_{k})\mid x^{-(r{+}1)}\,\psi_{p}(f)$ where $\psi_{p}$ is reduction modulo $p$. Then we obtain the better bound for $\phi(k)=\deg(\Phi_{k})\leq\deg(f)-(r{+}s{+}2)$ in Step 5 of Algorithm 4 (FindCycloFactors). If $f$ is palindromic then we have $r=s$.
Remark 4.6.
We use a simple technique for generating “random rationals”. We regard the rationals in ${\mathbb{Q}}_{>1}$ as being ordered lexicographically via the mapping $p/q\mapsto(p,q)$ with $\gcd(p,q)=1$. A new rational is chosen by jumping forward along this progression by a random amount; by always jumping forward, we avoid generating the same value twice.
4.2 Timings
We exhibit timings for the “difficult” case of many cyclotomic factors, since usually the first iteration whittles the list down to just those factors actually present, possibly with just very few false positives.
Each test polynomial is a product of distinct cyclotomic polynomials whose indexes come from a random subset of specified cardinality from the range $\{1,2,\ldots,R\}$. Each test set contains 10 such products. The table records the average degree, and the average time; no false positives were observed during testing.
Table 2: Timings for CyclotomicFactorIndexes
Index range Num Factors Avg degree Avg Time
$500$ $100$ 15 73115\,731 $0.10$
1 0001\,000 $50$ 15 69615\,696 $0.10$
1 0001\,000 $100$ 31 48731\,487 $0.22$
1 0001\,000 $200$ 61 44461\,444 $0.50$
We did observe a few false positives when using specially constructed square-free, palindromic polynomials such as $f=\prod_{k=2}^{201}(kx-1)(x-k)$ which have large “fixed divisors”; the fixed divisor is defined to be $\gcd\{f(n)\mid n\in{\mathbb{Z}}\}$. In all cases the false positives were indexes of cyclotomic polynomials of low degree (e.g. $3$, 44 and $6$).
5 Algorithms for Testing LRS-degeneracy
In this section we show that many LRS-degenerate polynomials exist, recall the algorithms presented in [9], and then present our new algorithm together with some sample timings.
5.1 Some properties of LRS-degenerate polynomials
We first show that infinitely many LRS-degenerate polynomials exist. Indeed, every member of the family of polynomials $\{x^{2}-n\mid n\in{\mathbb{Z}}_{\neq 0}\}$ is 2-LRS-degenerate; moreover, infinitely many of them are irreducible. In fact, all the polynomials in the family are LRS-equivalent to each other.
Directly from the definition, we can easily see that:
  • If $f\in{\mathbb{Z}}[x]$ is of the form $g(x^{k})$ for some $k\in{\mathbb{Z}}_{\geq 2}$ but not of the form $a_{d}\,x^{d}$ then $f$ is $\kappa$-LRS-degenerate for every $\kappa\geq 2$ dividing $k$ — we note that $f$ may also be LRS-degenerate for other orders too.
  • If $f\in{\mathbb{Z}}[x]$ is $k$-LRS-degenerate then so is any (non-zero) multiple of $f$.
  • Let $f\in{\mathbb{Z}}[x]$ be divisible by $x^{d}$ for some exponent $d\in{\mathbb{N}}$ then $f$ is $k$-LRS-degenerate iff $f/x^{d}$ is.
We observe that, for a given $f\in{\mathbb{Z}}[x]$ not of the form $a_{d}\,x^{d}$, it is computationally quick and simple to determine the largest $d,k\in{\mathbb{N}}$ such that $f(x)=x^{d}\,g(x^{k})$. If $d$ and $k$ are both even then $f$ is an even polynomial (i.e. f(x)=f(x)f(x)=f(-x)); and if $d$ is odd and $k$ even then $f$ is an odd polynomial (i.e. f(x)=f(x)f(x)=-f(-x)). We see that in both cases they are $2$-LRS-degenerate; of course, this is also immediate from the definitions of evenness and oddness.
There are many further examples. We mention two further quadratic polynomials: $x^{2}+3x+3$ is 6-LRS-degenerate, and $x^{2}+x+1=\Phi_{3}(x)$ is 3-LRS-degenerate. Lemma 5.1 immediately below summarises the LRS-degeneracy of cyclotomic polynomials. We can generate more LRS-degenerate polynomials as follows: let $f$ be LRS-degenerate and $g$ be not of the form $a_{d}\,x^{d}$ then $\mathop{\rm res}_{y}(f(y),\,g(xy))$ is LRS-degenerate since its roots are $\{\alpha/\beta\mid f(\alpha)=0,\;g(\beta)=0\}$; indeed, if $f$ and $g$ are irreducible then the resultant is likely irreducible too (provided $g$ remains irreducible over the extension generated by a root of $f$).
Lemma 5.1.
(cyclotomics are LRS-degenerate)
  • (a)
    For odd $k\in{\mathbb{N}}_{\geq 3}$ the polynomial $\Phi_{k}(x)$ is $\kappa$-LRS-degenerate for every factor $\kappa$ of $k$ (excluding $\kappa=1$).
  • (b)
    For even $k\in{\mathbb{N}}_{\geq 4}$ the polynomial $\Phi_{k}(x)$ is $\kappa$-LRS-degenerate for every factor $\kappa$ of $k/2$ (excluding $\kappa=1$).
Proof.
There is a 1–1 correspondence between primitive $k$-th roots of unity and $S_{k}=\{n\in{\mathbb{N}}\mid 1\leq n<k\text{ and }\gcd(n,k)=1\}$. Namely, the set of all primitive $k$-th roots is just $\{\zeta^{n}\mid n\in S_{k}\}$ where $\zeta$ is any primitive $k$-th root.
Let $\kappa>1$ be a factor of $k$, and let $\lambda=k/\kappa$. Then $\zeta^{j+\lambda}/\zeta^{j}$ is a primitive $\kappa$-th root of unity. We show how to find $j$ such that both $j$ and $j+\lambda$ are coprime to $k$, i.e. so that both $\zeta^{j}$ and $\zeta^{j+\lambda}$ are primitive $k$-th roots of unity.
Let the factorization of $k$ be $\prod_{j=1}^{m}p_{j}^{e_{j}}$ where the primes $p_{j}$ are distinct, and each $e_{j}>0$. We now construct a pair of suitable exponents $(i,j)$.
Case $k$ is odd. For each index $j$, pick non-zero residues $r_{j}\mod p_{j}$ such that $r_{j}{+}\lambda\not\equiv 0\pmod{p_{j}}$; note that all $p_{j}>2$, so a suitable $r_{j}$ exists. Using CRT construct $j\in\{1,2,\ldots,k-1\}$ such that $j\equiv r_{j}\pmod{p_{j}}$. Then $\gcd(j,k)=1$ and $\gcd(j{+}\lambda,k)=1$. Thus $\zeta^{j}$ and $\zeta^{j+\lambda}$ are primitive $k$-th roots, while $\zeta^{\lambda}=\zeta^{j+\lambda}/\zeta^{j}$ is a primitive $\kappa$-th root.
Case $k$ is even. Essentially the same argument works, but we must ensure that $\lambda$ is even; equivalently we must require that $\kappa$ divides $k/2$. ∎
Example 5.2.
The property that if $f$ is $k$-LRS-degenerate then it is $\kappa$-LRS-degenerate for all factors $\kappa>1$ dividing $k$ does not hold in general. For example, $x^{4}+2x^{2}+4x+2$ is $k$-LRS-degenerate only for $k=8$; and $x^{6}+3x^{5}+6x^{4}+6x^{3}+3$ is $k$-LRS-degenerate only for $k=18$.
5.2 Preprocessing prior to testing LRS-degeneracy
To simplify later discussions we shall assume that polynomials to be tested for LRS-degeneracy have been preprocessed. In [9] they describe some natural, simple preprocessing steps which we recall and expand here:
  • we may assume that $f(0)\neq 0$: just divide by appropriate power of $x$
  • we may assume that $f$ is non-constant and content-free in ${\mathbb{Z}}[x]$
  • we may assume that $f$ is square-free: just replace $f\leftarrow f/\gcd(f,f^{\prime})$
  • we may assume that $f$ is not of the form $g(x^{r})$ for some $r>1$:
    if $f(x)=g(x^{r})$ then $f$ is $k$-LRS-degenerate for each factor $k>1$ of $r$; if additionally $g(x)$ is $\kappa$-LRS-degenerate then $f$ is $r\kappa$-LRS-degenerate; $f$ may also be $k$-LRS-degenerate for other factors $k\mid r\kappa$.
  • Optionally, we may also simplify the coefficients using Algorithm 5 (ReduceCoefficients) from Section 5.2.1.
If we are interested in knowing only whether $f$ is LRS-degenerate, without actually determining the orders, then we can apply the following additional steps:
  • we may assume that $\gcd(f(x),f(-x))=1$: otherwise $f$ is clearly $2$-LRS-degenerate
  • we may assume that $f$ has no cyclotomic factors (with index >2>2): otherwise it is trivially LRS-degenerate — e.g. use the method of Section 4
  • we may also divide out any linear factors from $f$.
5.2.1 Preprocessing: coefficient reduction
We can use LRS-degeneracy equivalence to preprocess polynomials to be tested for LRS-degeneracy: the idea is to look for $\lambda$ and $\mu$ values which “simplify” the coefficients. A fully general approach appears to be potentially costly, probably entailing integer factorization. However, we can make a quick search for suitable $\lambda,\mu$; if we are lucky, we can simplify the polynomial.
Algorithm 5 (ReduceCoefficients) is a simple preprocessing method which produces an LRS-degeneracy equivalent polynomial whose coefficient size may be smaller than that of the input $f$. We run this algorithm twice: apply once to $f$ to obtain $f_{1}$, then apply again to $\mathop{\rm rev}(f_{1})$ to obtain $f_{2}$. The fully simplified polynomial is then $\mathop{\rm rev}(f_{2})$.
Remark 5.3.
If the input is $f=a_{d}x^{d}+a_{0}$, with $a_{d}\,a_{0}\neq 0$, then we can directly simplify to $x^{d}+1$. As presented, the algorithm will fail to fully simplify $f$ if it is unable to factorize $a_{0}$ or $a_{d}$.
Example 5.4.
Let $f=16x^{4}+80x^{3}+300x^{2}+1000x+3125$. The first run of the algorithm gives $\lambda=1/2$ and $f_{1}=x^{4}+10x^{3}+75x^{2}+500x+3125$. The second run, on input $\mathop{\rm rev}(f_{1})$, gives $\lambda=1/5$ and $f_{2}=5x^{4}+4x^{3}+3x^{2}+2x+1$. So the final simplified form of $f$ is $\mathop{\rm rev}(f_{2})=x^{4}+2x^{3}+3x^{2}+4x+5$ and the combined $\lambda$ is $5/2$.
Algorithm 5 (ReduceCoefficients)
1:Content-free polynomial $f=\sum_{j=0}^{d}a_{j}x^{j}\in{\mathbb{Z}}[x]$ with $d>0$ and $a_{0},a_{d}\neq 0$
2:
3:if $f(x)=g(x^{r})$ for some $r>1$ then apply this algorithm to $g(x)$ producing $\tilde{g}(x)$; return $\tilde{g}(x^{r})$ end if
[…]
5:if $g=1$ then return $f$ end if
[…]
7:for each “small” prime factor $p$ of $g$ do
8:  $m\leftarrow\min\{\lfloor\mu_{p}(a_{j})/j\rfloor\mid j=1,2,\ldots,d\}$
[…]
10:  $\lambda\leftarrow\lambda/p^{m}$
11:end for each
12:return $f(\lambda x)$ $\triangleright$ element of ${\mathbb{Z}}[x]$
5.3 Factorization and LRS-degeneracy
It is clear from the definition that when testing a polynomial $f$ for LRS-degeneracy we can work with the radical, $\mathop{\rm rad}(f)$, which may have lower degree. However we cannot, in general, use a finer factorization. We present an easy lemma, and then several examples which show that we cannot determine whether $f$ is LRS-degenerate by working independently on its irreducible factors.
Lemma 5.5.
Let $g(x)\in{\mathbb{C}}[x]$ be not LRS-degenerate then h(x)=g(x)g(x)h(x)=g(x)\,g(-x) is $k$-LRS-degenerate only for $k=2$.
Proof.
We show the contrapositive.
Suppose that $h(x)$ is $k$-LRS-degenerate for some $k\neq 2$. Then $h(x)$ has roots $\alpha,\beta\in{\mathbb{C}}$ such that $\alpha/\beta=\zeta_{k}$. WLOG $g(\alpha)=0$.
If $g(\beta)=0$ then $g$ is $k$-LRS-degenerate. Otherwise we have $g(-\beta)=0$; so $g(x)$ has a root pair with ratio $\alpha/\beta=-\zeta_{k}$, and hence $g(x)$ is LRS-degenerate (with order either $k/2$ or $2k$). ∎
Remark 5.6.
It is easy to see that if f(x)g(x)f(x)\,g(x) is LRS-degenerate then so is f(x)g(x)f(x)\,g(-x). Note that f(x)g(x)f(-x)\,g(x) is LRS-degeneracy equivalent to f(x)g(x)f(x)\,g(-x), and f(x)g(x)f(-x)\,g(-x) is LRS-degeneracy equivalent to f(x)g(x)f(x)\,g(x).
Example 5.7.
Here we exhibit some products of non-LRS-degenerate polynomials which are $k$-LRS-degenerate for $k\neq 2$. Let $f_{1}=5x^{2}+6x+5$ and $f_{2}=5x^{2}+8x+5$. Then neither $f_{1}$ nor $f_{2}$ is LRS-degenerate, but the product $f_{1}f_{2}$ is $4$-LRS-degenerate.
Let $g_{1}=7x^{2}+2x+7$, $g_{2}=7x^{2}+11x+7$ and $g_{3}=7x^{2}+13x+7$. None of these polynomials is LRS-degenerate, but the product of any pair is: $g_{1}g_{2}$ is $3$-LRS-degenerate, while $g_{1}g_{3}$ and $g_{2}g_{3}$ are $6$-LRS-degenerate.
We can generate more examples of degree 2 by applying Graeffe transforms: e.g. $G_{5}(f_{1})$ and $G_{5}(f_{2})$ are not LRS-degenerate but their product is 4-LRS-degenerate. This works for any $G_{k}$ with $k$ coprime to 4. Analogously, if we use the pair $g_{1},g_{2}$ then we can apply $G_{k}$ with $k$ coprime to 3; for the pairs $g_{1},g_{3}$ and $g_{2},g_{3}$ we need $k$ coprime to $6$.
We can generate similar examples with higher degrees, for instance: let $h$ be an irreducible polynomial which is not LRS-degenerate, and compute $h_{i}=\mathop{\rm res}_{y}(f_{i}(y),h(xy))$ for $i=1,2$. Most likely $h_{1}$ and $h_{2}$ are irreducible and not LRS-degenerate, but their product is $4$-LRS-degenerate.
Example 5.8.
We can generate more pairs of non-LRS-degenerate polynomials whose product is LRS-degenerate as follows. Pick an index $k\geq 7$, and a random polynomial $g\in{\mathbb{Z}}[x]$ with $\deg(g)<\phi(k)$. Compute $m(y)=\mathop{\rm res}_{x}(\Phi_{k}(x),\,y-g(x))$; if this is reducible, pick a different $g$. Let $S$ be the set of irreducible factors of $\mathop{\rm res}_{y}(m(y),\Phi_{k}(xy))$; then in many cases the factors are not LRS-degenerate, but the product of any pair is $\kappa$-LRS-degenerate for some $\kappa>1$ dividing $k$.
5.4 Recalling CDM Algorithm 1
We recall briefly the two algorithms for detecting LRS-degenerate polynomials which were published in [9]. The presentation in that paper was mostly concerned with determining solely whether a given polynomial is LRS-degenerate, rather than determining the LRS-degeneracy orders.
These algorithms assume that the input $f\in{\mathbb{Z}}[x]$ has been preprocessed (see Section 5.2) so that it has degree at least $2$, and is content-free, square-free and $f(0)\neq 0$.
The first algorithm in [9] computes $R_{f}=\mathop{\rm res}_{y}(f(y),f(xy))/(x-1)^{d}$ where $d=\deg(f)$. So $R_{f}$ is a palindromic polynomial of degree $d^{2}-d$, since we assume that $f$ is square-free. It then determines whether $R_{f}$ has any cyclotomic factors (with index $\geq 2$), since $f$ is $k$-LRS-degenerate iff $\Phi_{k}(x)$ divides $R_{f}$. In practice, they observe that this method becomes rather slow for $\deg(f)>25$ because the resultant computation is costly. Also they sought cyclotomic factors by first computing a factorization into irreducibles, which can be slow in some cases. Our Algorithm 4 (FindCycloFactors) or that from [14] would make the detection of cyclotomic factors much faster, but would not reduce the cost of computing the resultant. Our Algorithm 4 (FindCycloFactors) produces also the orders of LRS-degeneracy.
5.5 Recalling CDM Algorithm 2
The second algorithm in [9] replaces the costly resultant computation by a succession of simpler resultant computations. Effectively it searches for the lowest order of LRS-degeneracy. They report that it is usefully faster than their first algorithm if the polynomial is indeed $k$-LRS-degenerate for some small $k$ — but if it is not LRS-degenerate then their first algorithm is often faster.
Recall that we assume that $\gcd(f(x),f(-x))$ is constant, so $f$ is not 2-LRS-degenerate. They try all candidate orders $k=3,4,\ldots,5d^{2}$: for each candidate compute $R_{f,k}=\mathop{\rm res}_{y}(f(y),y^{k}-x)$, and test whether $R_{f,k}$ is square-free; if not then $f$ is $k$-LRS-degenerate. We note that $R_{f,k}=G_{k}(f)$, the $k$-th Graeffe transform of $f$.
Remark 5.9.
For some indexes $k$ we have that $\deg(\Phi_{k})\ll k$, so we can consider computing $R^{*}_{f,k}=\mathop{\rm res}_{y}(f(xy),\Phi_{k}(y))$ and checking whether it is square-free. For such $k$ the advantage is that $R^{*}_{f,k}$ has rather lower degree than $R_{f,k}$. In practice, the cost of computing $R^{*}_{f,k}$ was sometimes higher than the cost of computing $R_{f,k}$ because $y^{k}-x$ has a simple, sparse structure whereas $\Phi_{k}(y)$ is not sparse.
Remark 5.10.
The upper bound for the loop, $5d^{2}$, is actually a bound on the inverse of Euler’s totient function for the degrees they were able to handle (see [7]). The basis for Algorithm 1 of [9] tells us that if $f$ is $k$-LRS-degenerate then $\Phi_{k}\mid R_{f}$, thus by considering degrees, $\phi(k)\leq d^{2}-d$; whence the upper bound (for $d\leq 75$). As we already mentioned, sequence A355667 at OEIS [16] implies a dynamic bound, which is needed when $d>75$. Our Section 5.9.1 introduces a conjecture which, if true, would let us use a much lower bound.
5.6 The modular algorithm
We now present our modular approach whose existence was hinted at in [9]. Our approach employs “sophisticated brute force”, and comprises two parts: one testing whether a polynomial is $k$-LRS-degenerate for a specific order $k$, the other choosing which orders $k$ to test. Recall that the input is content-free, square-free $f\in{\mathbb{Z}}[x]$ with $f(0){\neq}0$, degree $\geq 2$ and $\gcd(f(x),f(-x)){=}1$. The following subsections give further details and justifications for our new Algorithm 6 (LRSDegeneracyOrders).
Remark 5.11.
In Algorithm 6 (LRSDegeneracyOrders) the loop from Step 6 to Step 13 performs 33 iterations because that gave a good compromise between speed of computation and exclusion of false positives.
5.7 The modular algorithm: supporting lemma
Lemma 5.12.
Let $f\in{\mathbb{Z}}[x]$ be $k$-LRS-degenerate for some order $k\in{\mathbb{N}}_{\geq 2}$. Let ${\mathbb{F}}_{q}$ be a finite field containing a primitive $k$-th root of unity, $\zeta_{k}$. Let $\psi:{\mathbb{Z}}\to{\mathbb{F}}_{q}$ be the canonical ring homomorphism, with natural extension to $\psi:{\mathbb{Z}}[x]\to{\mathbb{F}}_{q}[x]$. Let $g=\psi(f)$, and suppose that $\deg(g)=\deg(f)$ then $\deg\bigl(\gcd(g(x),g(\zeta_{k}x))\bigr)\geq 1$ in ${\mathbb{F}}_{q}[x]$.
Proof.
Let ${\mathbb{Z}}[\xi_{k}]$ be the ring extension of $\mathbb{Z}$} by a $k$-th root of unity, $\xi_{k}$. We have a unique ring homomorphism $\theta:{\mathbb{Z}}[\xi_{k}]\to{\mathbb{F}}_{q}$ sending $\xi_{k}$ to $\zeta_{k}$, which extends naturally to $\theta:{\mathbb{Z}}[\xi_{k}][x]\to{\mathbb{F}}_{q}[x]$; observe that $\theta|_{{\mathbb{Z}}[x]}=\psi$. Regarding ${\mathbb{Z}}[x]$ as embedded in the ring ${\mathbb{Z}}[\xi_{k}][x]$, let $h=\gcd(f(x),f(\xi_{k}x))$, so by assumption $\deg(h)\geq 1$. Since $\deg(g)=\deg(f)$, we have also $\deg(\theta(h))=\deg(h)$. By construction $h(x)\mid f(x)$ and $h(x)\mid f(\xi_{k}x)$, so $\theta(h(x))\mid\theta(f(x))$ and $\theta(h(x))\mid\theta(f(\xi_{k}x))$. Now $\theta(f(x))=g(x)$ and $\theta(f(\xi_{k}x))=g(\zeta_{k}x)$; thus $\theta(h(x))\mid\gcd(g(x),g(\zeta_{k}x))$. ∎
We are especially interested in the case where $q=1+rk$ is prime. In this case, all primitive $k$-th roots of unity are elements of ${\mathbb{F}}_{q}$. Let $Z\subset{\mathbb{F}}_{q}$ be the set of these primitive roots. By the lemma, for each $\zeta_{k}\in Z$ we have $\deg\bigl(\gcd(g(x),\,g(\zeta_{k}x))\bigr)\geq 1$ in ${\mathbb{F}}_{q}[x]$.
5.8 Test for $k$-LRS-degeneracy
Here we present our sub-algorithm for testing whether $f$ is $k$-LRS-degenerate: this corresponds to the loop controlled by Step 5. We assume that $k>2$ since 2-LRS-degeneracy is easy to test. Our test is one-sided: it may produce a false positive, namely, report that $f$ is $k$-LRS-degenerate when it is actually not. However, if $f$ is not $k$-LRS-degenerate, this is likely to be detected quickly.
A candidate order $k$ can be fully verified by checking that $\mathop{\rm res}_{y}(f(y),\Phi_{k}(xy))$ is not square-free; we did not put this verification into the algorithm because it can be quite costly when $k$ is large. In Section 5.5 when discussing the second algorithm from [9], we observed that computing $R_{f,k}=\mathop{\rm res}_{y}(f(y),\,y^{k}-x)$ is typically faster than computing $\mathop{\rm res}_{y}(f(y),\,\Phi_{k}(xy))$; we cannot take this “short-cut” here because the non-square-freeness of $R_{f,k}$ indicates only that $f$ is $\kappa$-LRS-degenerate for some factor $\kappa$ of $k$.
Algorithm 6 (LRSDegeneracyOrders)
1:Polynomial $f\in{\mathbb{Z}}[x]$, square-free with $f(0)\neq 0$
2:
[…]
5:for each candidate order $k\in K$ do
6:  for $i=1,2,3$ do
7:   Pick a prime $p$ with $p\equiv 1\pmod{k}$
8:   Compute $\zeta_{k}$, a primitive $k$-th root of unity in ${\mathbb{F}}_{p}$
9:   for $j=1,2,\ldots,\lfloor k/2\rfloor$ do
10:     if $\gcd(j,k)=1$ and $\gcd(\bar{f}(x),\bar{f}(\zeta_{k}^{j}x))=1$ then
[…]
12:   end for
13:  end for
14:  Append $k$ to $L$ as a probable LRS-degeneracy order
15:end for each
16:return $L$
In Step 7 we pick a suitable finite field and employ Lemma 5.12. By Dirichlet’s theorem we know that there are infinitely many primes of the form $1+ks$. To lower the risk of false positives we choose a prime $p>8(\deg f)^{2}$ and such that $s\geq 64$; we avoid small $s$ so that the vast majority of field elements are not $k$-th roots of unity, and we want all ratios of roots of $f$ to cover only a small proportion of the field elements. We obtain a primitive $k$-th root of unity via the hint in Example 2.2. Our implementation chooses the primes randomly, so it is harder to construct “pathological” inputs.
In the loop at Step 9 we can stop iterating at $\lfloor k/2\rfloor$ because for $j>k/2$ we have $\zeta_{k}^{j}=1/\zeta_{k}^{k-j}$, so letting $g=\gcd(\bar{f}(x),\bar{f}(\zeta_{k}^{k-j}x))$ we deduce that $\gcd(\bar{f}(x),\bar{f}(\zeta_{k}^{j}x))=u\,g(\zeta_{k}^{j}x)$ for some unit $u$. So both gcds have the same degree. Here we write $\bar{f}$ to denote the canonical image of $f$ in ${\mathbb{F}}_{p}[x]$.
To further reduce the chance of a false positive, we perform the analogous test for up to two more primes: the loop at Step 6 performs at most 3 iterations. If all these checks pass then “very likely” $f$ is indeed $k$-LRS-degenerate, so we return a positive result. Quantifying the probability of a false positive would require assuming a distribution on the input polynomials, but there is no obvious, reasonable choice.
5.9 Which $k$ values to try?
In Step 3 of Algorithm 6 (LRSDegeneracyOrders) we need an initial set of candidate orders to test. Recall that we assume that $f$ is square-free in ${\mathbb{Z}}[x]$; we write $d=\deg(f)$. One possible set comprises all $k$ such that $\phi(k)\leq d^{2}-d$, because if $f$ is $k$-LRS-degenerate then $\Phi_{k}(x)$ divides $\mathop{\rm res}_{y}(f(xy),f(y))/(x-1)^{d}$ which has degree $d^{2}-d$. Below we explain how to reduce this set somewhat.
We can refine the set of $k$ values to try as follows. Assume that $f$ is $k$-LRS-degenerate, and let $\alpha,\beta$ be roots of $f$ for which $\alpha/\beta=\zeta_{k}$. We consider two cases:
  • If there is an irreducible factor $g$ of $f$ having both $\alpha$ and $\beta$ as roots then the algebraic extension $L={\mathbb{Q}}(\alpha,\beta)$ has degree $d_{\alpha}d_{\beta}$ where $d_{\alpha}=\deg(g)$ and $d_{\beta}<d_{\alpha}$. Since $L$ contains ${\mathbb{Q}}(\zeta_{k})$ we have that $\phi(k)\mid d_{\alpha}d_{\beta}$.
  • Otherwise there are distinct irreducible factors $g_{\alpha}$ and $g_{\beta}$ being the minimal polynomials for $\alpha$ and $\beta$ respectively. Let $d_{\alpha}=\deg(g_{\alpha})$ and $d_{\beta}=\deg(g_{\beta})$; and we may assume that $d_{\beta}\leq d_{\alpha}$. Clearly $d_{\alpha}+d_{\beta}\leq d$. As before we have that $\phi(k)\mid d_{\alpha}d_{\beta}$
[…]
Algorithm 7 (LRSOrderCandidates)
[…]
2:
[…]
4:for $\alpha=1,2,\ldots,d$ do
5:  for $\beta=1,2,\ldots,\alpha-1$ do
6:   if $\alpha\beta$ is even then append $\alpha\beta$ to $D$ end if
7:  end for
8:end for
9:for $\alpha=2,4,6,\ldots,d/2$ do
10:  Append $\alpha^{2}$ to $D$
11:end for
12:return list of all $k>2$ up to $\phi^{-1}_{max}(d^{2}-d)$ such that $\phi(k)$ divides an element of $D$
5.9.1 Conjectured short-cut
If we are just interested in the smallest order $k$ such that $f$ is $k$-LRS-degenerate then, dependent on a conjecture, we can reduce the list of $k$ to try considerably.
  • We conjecture that if $f$ is LRS-degenerate then the minimal order $k$ satisfies $\phi(k)\leq\deg(f)$.
  • Furthermore, we conjecture that if $f$ is irreducible and LRS-degenerate then it is $k$-LRS-degenerate for some order $k$ with $\phi(k)\mid\deg(f)$.
Example 5.13.
Let $f=\Phi_{3}\Phi_{5}$, so $\deg(f)=6$. Then $f$ is obviously $k$-LRS-degenerate for $k=3$ and $k=5$. In fact, $f$ is also $15$-LRS-degenerate, yet $\phi(15)=8>\deg(f)$. So if we want to find all orders $k$ then the conjectured limit does not apply.
5.10 Testing for LRS-degeneracy via evaluation
We considered using the method of Section 4.1 to identify the cyclotomic factors in $R_{f}$. A key point is that, for any evaluation point $b$, we have $\mathop{\rm res}_{x}(f(x),f(bx))=R_{f}(b)$, where $R_{f}$ is the resultant from Section 5.4. Observe that the numerical resultant $\mathop{\rm res}_{x}(f(x),f(bx))$ can be computed directly and quickly without having to compute explicitly $R_{f}(y)$ as a polynomial. We can also use the algorithm of Section 5.9 to supply an initial list of candidate indexes.
We tried this implementation, but it was slower than our modular method of Section 5.6: the cost of computing the numerical resultants (in CoCoALib) was too high.
5.11 Timings
In Table 3 we present some timings of our implementation of Algorithm 6 (LRSDegeneracyOrders). We tried the three largest examples from [9], each of degree 24; but each example took less than 0.1s (to compute all LRS-degeneracy orders). Instead we generated some higher degree polynomials: four with uniform random integer coefficients from $2^{10}$ to $2^{10}$ (a larger coefficient range made essentially no difference); and three products of cyclotomic polynomials which have many orders. As expected, none of the random polynomials was LRS-degenerate.
Table 3: Timings for LRSDegeneracyOrders
Description Deg Unverified Verified
Random $25$ $0.02$ $0.02$
Random $50$ $0.03$ $0.03$
Random $100$ $0.19$ $0.20$
Random $200$ $1.7$ $1.7$
$\Phi_{51}\Phi_{65}\Phi_{77}$ $140$ $0.8$ $35$
$\Phi_{165}\Phi_{183}$ $200$ $2.2$ $118$
$\Phi_{123}\Phi_{185}\Phi_{217}\Phi_{299}\Phi_{319}\Phi_{323}$ 1 2361\,236 3 2433\,243
In none of our tests did we observe any false positive. Also, for polynomials with larger LRS-degeneracy orders, the costs of verification dominate: for instance, we did not perform verification for the degree 1 2361\,236 polynomial as fifteen of the candidate orders are at least 22 75522\,755, and the largest order is 103 037103\,037.
Remark 5.14.
For the verification we checked that $\mathop{\rm res}_{y}(f(xy),\Phi_{k}(y))$ has a repeated factor; this is essentially the technique used inside the loop in CDM Algorithm 2 — see Section 5.5. To lower the computational cost we used the fact that the resultant is “multiplicative”, namely $\mathop{\rm res}(f,g_{1}g_{2})=\mathop{\rm res}(f,g_{1})\,\mathop{\rm res}(f,g_{2})$, and the classical product from Lemma 2.7 (a). We do this because it is decidedly quicker to compute $\mathop{\rm res}_{y}(f(xy),y^{k}-1)$ rather than $\mathop{\rm res}_{y}(f(xy),\Phi_{k}(y))$.
6 Conclusion
We have presented new, practical algorithms for:
  • testing whether a polynomial in ${\mathbb{Z}}[x]$ is cyclotomic, and if so, for finding its index.
  • finding the indexes of all cyclotomic factors of a given polynomial in ${\mathbb{Z}}[x]$ with a low likelihood of false positives.
  • testing whether a polynomial in ${\mathbb{Z}}[x]$ is LRS-degenerate, and if so, determining all LRS-degeneracy orders, with a low likelihood of false positives. Currently, final verification of the candidate orders dominates the computation time.
The algorithms presented here are implemented as part of in CoCoALib [1] (from version 0.99856). The software is free and open source; it is available from https://cocoa.dima.unige.it/